Suppose that there are only two goods produced in the economy, Call center services and banking services. Prices (P), quantities (Q) and the number of workers (W) occupied in the production of each good for year 1 and year 2 are given by the following table : Year 1 Year 2 Two products P1 Q1 W1 P2 Q2 W2 Call centre 10 100 50 12 100 50 Banking 10 200 50 12 230 60

 Using year 1 prices, what is real GDP in year 2? What is the growth rate of real GDP?

What is the rate of inflation using the GDP deflator?

 Using year 1 prices, what is real GDP per worker in year 1 and year 2? What is labour productivity growth between year 1 and year 2 for the whole economy?


1
Expert's answer
2021-12-02T20:06:11-0500


(a)

Nominal GDP for year 1

Call center services:

"=10\times 100=1,000"

Banking services:

"=10\times 200=2,000"

Total"=3,000"

Nominal GDP for year 2

Call center services:

"=12\times 100=1,200"

Banking services:

"=12\times 230=2,760"

Total"=3,960"


(b)

Real GDP:

Given by multiplying base year prices by current year quantities.

Base year= year 1.

Call center services:

"=10\times 100=1000"

Banking services:

"=10\times 230=2300"

Total"=3300"


(c)

Rate of inflation:

"GDP=\frac{Nominal GDP}{Real GDP}\times100"

"=\frac{3960}{3300}\times100=120"


(d)

Real GDP per worker: given by dividing GDP at constant prices by size of population

Year 1:

"=\frac{3000}{100}=30"

Year 2:

"=\frac{3960}{110}=36"